OpenStudy (anonymous):

what is the distance between the origin of a cartesia coordinate system and the point (4,-6)

OpenStudy (anteater):

If you use the distance formula, \[d = \sqrt{(y _{2}- y _{1})^2 + (x _{2}-x _{1})^2}\]

OpenStudy (anteater):

And they give you the coordinates of one point (x,y) = (4, -6) and tell you that the other point is the origin. What are the coordinates of the origin?

OpenStudy (anteater):

What is meant by the "origin" is the point (0,0), the point where the x and y axes intersect.

OpenStudy (anteater):

So, you have two points: \[(x _{1},y _{1}) = (0,0) and (x _{2},y _{2}) = (4, -6)\] Using that information and the distance formula, what answer do you get?

OpenStudy (anonymous):

i dont know

OpenStudy (anteater):

If you plug in the information: \[d = \sqrt{(x _{2}-x _{1.)^2}+(y _{2}-y _{1})^2} = \sqrt{(4 - 0)^2 + (-6-0)^2}\]

OpenStudy (anteater):

When you do the arithmetic, what answer do you end up with?